OpenStudy (anonymous):

find the minimum velocity to be imparted to a body so that it escapes from solar system

OpenStudy (anonymous):

Its kinetic energy must exceed the potential energy due to gravitation from the sun, (1/2) m v^2 > G m M / r where G is universal gravitation constant, m and M are its mass and solar mass and r is how far from the sun it starts.

OpenStudy (anonymous):

to escapes from Earth k.e energies should exceed the total energy (1/2) m v^2 > ke + pe now pe = -gmM/r (r= radius of earth + height above the surface of earth ) ke= (1/2) m v^2 but v^2 =gm/r (r= radius of earth + height above the surface of earth ) KE = 1/2gmM/r therefore T .E > ke+ pe T .E > -gmM/2r (r= radius of earth + height above the surface of earth ) (1/2) m v^2 > -gmM/2r (r= radius of earth + height above the surface of earth )

OpenStudy (anonymous):

The escape velocity needed will depend on where the object starts and which gravitational field / potential is dominant. I assumed "Sun" from "solar system," but perhaps "Earth" is what is wanted. Thus, mass of earth and distance from it or mass of sun and distance from it would prevail.